Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Evaporator_2  Not logged in ELOG logo
Message ID: 166     Entry time: Thursday, October 09, 2014, 17:41
Author: John Wong 
Target Material: Nb 
Target Oven W.O#:  
Source: TARGET BOAT 
Status: Done 
Subject: Conditioning Nb targets (for Bob)  

Thursday, October 09, 2014, 17:36

Target boat consists of 5 Nb discs.  Each disc is wrapped with Nb foil, and to be heated to 1400C for 4 hours.

Chamber has been pumped down to 6.5e-7 Torr.  HS valves open - no leaks detected.

TGHT has ramped to 330A to maintain the 1400C temperature. 

At 21:30, the target will be cooled down.

Temperature can be monitored from this link:

https://www.dropbox.com/sh/kor6gnutnlimqxw/AAAee2yb-Hj3JEbBysrAW2oYa?dl=0


 Thursday, October 09, 2014, 22:48

Samples have been heated at 1420C for over 4 hours.  

Pressure at 6.6 e-7 Torr. 

Cooling process initiated, ramping down at 2A/min. 


 Friday, October 10, 2014, 13:50

Samples were successfully conditions; no signs of any damage.


 

Attachment 1: samples@1400C.png  429 kB  Uploaded Thursday, October 09, 2014, 23:53  | Show | Hide all
ELOG V2.9.2-2455